# MorphoSeg: An Uncertainty-Aware Deep Learning Method for Biomedical Segmentation of Complex Cellular Morphologies

### v1

@@ Abstract
Deep learning has revolutionized medical and biological imaging, particularly in segmentation tasks. However, segmenting biological cells remains challenging due to the high variability and complexity of cell shapes. Addressing this challenge requires high-quality datasets that accurately represent the diverse morphologies found in biological cells. Existing cell segmentation datasets are often limited by their focus on regular and uniform shapes. In this paper, we introduce a novel benchmark dataset of Ntera-2 (NT2) cells, a pluripotent carcinoma cell line, exhibiting diverse morphologies across multiple stages of differentiation, capturing the intricate and heterogeneous cellular structures that complicate segmentation tasks. To address these challenges, we propose an uncertainty-aware deep learning framework for complex cellular morphology segmentation (MorphoSeg) by incorporating sampling of virtual outliers from low-likelihood regions during training. Our comprehensive experimental evaluations against state-of-the-art baselines demonstrate that MorphoSeg significantly enhances segmentation accuracy, achieving up to a 7.74% increase in the Dice Similarity Coefficient (DSC) and a 28.36% reduction in the Hausdorff Distance. These findings highlight the effectiveness of our dataset and methodology in advancing cell segmentation capabilities, especially for complex and variable cell morphologies. The dataset and source code is publicly available at https://github.com/RanchoGoose/MorphoSeg.
@@ 1 Introduction
To address the aforementioned challenges, we propose a novel unknown-aware enhancement algorithm based on the SOTA medical segmentation framework TransUNet [1]. Our approach involves sampling virtual outliers from the low-likelihood regions of the class-conditional distributions during the training stage to improve the decision boundaries of the segmented objects. This strategy enhances the robustness of the deep learning models, making them effective solutions in such complex scenarios. By incorporating uncertainties into the training objective, our proposed deep learning model achieves strong segmentation of cells with random shapes. When compared against several benchmarks, we demonstrate significant improvements with our proposed dataset: the mean Dice Similarity Coefficient (DSC) increased from 80.35% to 86.57% and the mean Hausdorff Distance (HD95) decreased from 21.98% to 15.75%. These results confirm that our approach not only mitigates data scarcity but also enhances the model’s generalization capabilities across varied cell morphologies. We have provided the full code and pre-trained models with simple inference instructions to facilitate easy implementation and further research by field researchers.
@@
    Development of a pluripotent carcinoma cell repository, specifically curated to support the development and validation of advanced biological cell segmentation models.
@@
    Proposal of a practical and scalable framework for cell segmentation that leverages synthetic outliers and data augmentation to significantly enhance the performance of the SOTA cell segmentation deep-learning models, particularly in scenarios typical of limited data availability.
@@
    Comprehensive quantitative and qualitative analyses highlighting that the proposed model performs favourably compared to existing techniques.
@@ 2 Related Work
In recent years, deep learning techniques in computer vision have achieved promising performance in medical image segmentation. Encoder-decoder-based CNN structures have been widely applied, with U-Net and its variants significantly advancing the field [2, 13, 19, 20, 21]. Building upon this structure, UNet++ incorporates densely connected links to enhance the fusion of multi-scale features, improving the analysis of medical images across CT, MRI, and EM modalities [22]. Furthermore, nnU-Net, a self-configuring U-Net, has shown remarkable performance across various segmentation tasks with minimal manual parameter tuning [23, 24]. In terms of segmenting cell structures, the Cellpose model utilizes a two-headed output architecture to predict cell probabilities and spatial organization, effectively handling cells of regular shapes [6]. However, its performance is limited for irregularly shaped cells, as often encountered in advanced disease stages.
@@
Semantic segmentation is pivotal for clinical tasks such as disease detection, differential diagnosis, survival prediction, therapy planning, and treatment response assessment [25, 26, 4, 27, 28, 29]. Besides deep learning methods based on CNN structures, attention mechanisms [30] are extensively employed in semantic segmentation to prioritize salient features. The integration of Vision Transformers (ViTs) [31] into medical segmentation networks has enhanced performance by leveraging self-attention mechanisms [32, 33, 34]. Some research focuses on combining the strengths of transformers with traditional CNN architectures, creating hybrid encoder structures. TransUNet combines CNNs for spatial information extraction with transformers for capturing long-range dependencies, featuring a U-Net structure with self-attention mechanisms to ensure comprehensive image analysis [1]. This integration makes TransUNet particularly effective for detailed medical and biological image segmentation, where preserving spatial hierarchies is crucial. Other research includes employing channel attention to capture boundary-aware features for enhancing polyp segmentation [35]. CellViT replaces the CNN encoder with a transformer block in the U-Net architecture [36]. The recent SAMed model extends the capabilities of the Segment Anything Model (SAM) [37] to medical imaging by employing a Low-rank Adaptation (LoRA) fine-tuning strategy to adapt the large-scale image encoder for medical contexts effectively [3]. This method illustrates the advantage of using extensive pre-trained models for specialized medical segmentation tasks. However, SAM is designed for multi-class segmentation, posing limitations for binary segmentation scenarios.
@@ 2.2 Virtual Outliers Synthesis
Outliers in data mining and statistics are typically considered representations of abnormalities, discordants, deviants, or anomalies [38, 39]. In the context of deep learning, outliers are often viewed as anomalies compared to inliers, with multiple deep learning methods developed to detect these outliers [40, 39]. Unlike traditional approaches that focus on outlier detection, recent works in computer vision have explored synthesizing outliers to enhance the accuracy and robustness of deep learning models, particularly in scenarios where datasets are insufficient, such as the medical image classification [39].
Further research has leveraged synthetic data to quantify uncertainty, primarily for out-of-distribution (OOD) detection [41, 42, 4]. OOD detection aims to identify conditions unrepresented in the training dataset [43]. In the realm of OOD detection, the challenge is to detect OOD samples; in our case, these include unseen cells in annotations that are difficult for precise human annotation and cells at transitional stages that can be considered OOD. The subtle differences between foreground elements in medical images turn OOD detection into a near-OOD problem, where traditional methods often fall short [44, 45, 46, 47]. In our research, we treat segmentation targets, specifically NT2 cells at various stages, as OOD data. Virtual outlier synthesis has proven effective for OOD detection [41], inspiring us to adapt this approach to enhance the robustness and accuracy of our model for cell segmentation.
Synthesis methods for medical and biological image segmentation remain relatively underexplored. Few synthesis methods have been applied to enhance the training of deep learning models in this area [48, 49, 50]. Our proposed method dynamically estimates class-conditional Gaussian distributions during training, shaping the uncertainty surface over time with our specially formulated loss.
Figure 2: The overall framework of the proposed method, following the architecture of TransUNet [1]. The process for virtual outlier sampling is depicted with red dotted lines. In this framework, we model the feature representations as class-conditional Gaussians and sample virtual outliers $v$ from regions of low likelihood. These virtual outliers contribute to generating the uncertainty loss for regularization. The synthetic segmentation map, incorporating the uncertainty estimation branch $\mathcal{L}_{\text{uncertainty}}$, is trained concurrently with the mixed segmentation loss $\mathcal{L}_{\text{seg}}$ in Eq. (12).
@@ 3 Methodology
The overall framework of our approach is depicted in Figure 2. We utilize the TransUNet [1] architecture as the foundational framework for our method. Initially, high-resolution images are subdivided into small patches to align with the input constraints of the vision transformer and enhance the model’s ability to capture cellular details at the pixel level (details in Section 4.2). For the original TransUNet, there are two types of encoders for feature extraction: a hybrid encoder combining a pre-trained CNN and a pure Transformer encoder. We present the hybrid structure in Figure 2 as it provides better results (shown in Section 5.3). It is noteworthy that our method is scalable to different deep learning frameworks since the sampling method works on the raw segmentation predictions (details in Section 3.2). The rest of this section is orgnized as follows: Section 3.1 provides preliminary information and an overview of the TransUNet framework for image segmentation tasks. In Section 3.2, we present our robust cell segmentation enhancement via virtual outlier sampling. Section 3.3 introduces the overall uncertainty-aware training objective for MorphoSeg.
@@ 3.1 TransUNet for Cell Segmentation
Given an image $\bm{\mathrm{x}}\in\mathbb{R}^{H\times W\times C}$ with a spatial resolution of $H\times W$ and $C$ channels, our objective is to predict the corresponding pixel-wise label map of the same dimensions. Traditionally, this is achieved by training a CNN (e.g., U-Net) that encodes the image into high-level feature representations and subsequently decodes these representations back to the original spatial resolution. To enhance segmentation accuracy, particularly for cell segmentation, we incorporate a Vision Transformer with self-attention mechanisms for instance segmentation without losing fine-grained information.
Image Sequentialization. For a pure Transformer encoder structure, following [31], we first perform tokenization by reshaping the input $\bm{\mathrm{x}}$ into a sequence of flattened 2D patches $\bm{\mathrm{x}}^{i}_{p}\in\mathbb{R}^{P^{2}\cdot C}|i=1,..,N\}$, where each patch is of size $P\times P$ and $N=\frac{HW}{P^{2}}$ is the number of image patches (i.e., the input sequence length).
Patch Embedding. We map the vectorized patches $\bm{\mathrm{x}}_{p}$ into a latent $D$-dimensional embedding space using a trainable linear projection. To encode the patch spatial information, we learn specific position embeddings which are added to the patch embeddings to retain positional information as follows:
@@
The concatenated sequence $\mathbf{T}$ is processed through the Transformer encoder layers to yield the final representations. Following the the architecture of the vanilla ViT [31], we employ multi-headed self-attention (MSA) [30] (Eq. 3) and multi-layer perceptron (MLP) blocks (Eq. 4). Layer normalization (LN) is applied before each block (Eq. 5), as described in the following equations:
@@
Due to the limited dimensions of ViT input tokens, input image patch sizes $\frac{H}{P}\times\frac{W}{P}$ are often much smaller than the original image resolution $H\times W$ in high-resolution medical images, leading to a loss of low-level details (e.g., cell shape and boundary). TransUNet addresses this by using a CNN-Transformer hybrid model, where a CNN first extracts a feature map from the input. Patch embedding is then applied to $1\times 1$ patches from the CNN feature map instead of raw images, ensuring better retention of details. This method’s efficacy is demonstrated in the comparative analysis of different encoders presented in Table 1 in Section 5.3.
Cascaded Upsampler. Instead of naive upsampling in a ViT, a Cascaded Upsampler (CUP) is implemented, consisting of multiple upsampling steps to decode the hidden features and output the final segmentation mask. After reshaping the sequence of hidden features $\bm{\mathrm{z}}_{L}\in\mathbb{R}^{\frac{HW}{P^{2}}\times D}$ to the shape $\frac{H}{P}\times\frac{W}{P}\times D$, CUP is instantiated by cascading multiple upsampling blocks to reach the full resolution from $\frac{H}{P}\times\frac{W}{P}$ to $H\times W$. Each block consists of a 2×2$\times$2 × upsampling operator, a $3\times 3$ convolution layer, and a ReLU layer successively. CUP, together with the hybrid encoder, forms a U-shaped architecture that enables feature aggregation at different resolution levels via skip connections. The upsampling process is also illustrated in Figure 2.
@@ 3.2 Robust Cell Segmentation via Virtual Outlier Sampling
To enhance the robustness of our model and accurately simulate the different stages of pluripotent carcinoma cells, we propose the sampling of virtual outliers during the training process. Traditional deep learning models for image segmentation typically utilize CNN frameworks, which compress the latent space into lower dimensions through feature extraction. In contrast, our model is built on a Vision Transformer architecture, which poses challenges in synthesizing virtual outliers directly in the high-dimensional pixel space due to its complexity and the optimization difficulties involved. Despite the CNN-Transformer hybrid structure as its encoder in the form of TransUNet, the latent features of high-dimensional images remain in a large matrix. To address this, we opt to sample directly from the output logits of the segmentation map, denoted by $s(\mathbf{x}_{i})$, instead of sampling from the latent space.
@@
where $v_{k}\sim\mathcal{N}(\widehat{\mu}_{k},\widehat{\Sigma}_{k})$ represents the sampled virtual outliers for class $k$. These samples are chosen from a sublevel set defined by a small $\epsilon$, ensuring that the outliers are near the decision boundary of the class. We hypothesize that most of the cell pixels are near the boundary of existing cells as shown in Figure 3. Compared to the ground truth masks shown in the second subfigure in Figure 3, the synthesized outliers contain more information about the cells, which are difficult for humans to annotate due to their small size. This strategy aims to enhance the model’s sensitivity to subtle yet critical variations in cell morphology that might indicate different developmental stages or pathological conditions.
@@
3.3 Unknown-aware Training Objective
The objective is to develop an uncertainty-aware deep learning model capable of predicting a segmentation map $\hat{S}$ for input cell images. To this end, we introduce a new training objective for unknown-aware learning, which utilizes virtual outliers as discussed in Section 3.2. We start by defining the standard Dice loss between the predicted pixel scores $\hat{y}$ and the ground truth targets $y$:
$\mathcal{L}_{\text{Dice}}(f(\mathbf{x};\theta),y)=1-\frac{2\sum_{i}p_{i}g_{i}+\epsilon}{\sum_{i}p_{i}^{2}+\sum_{i}g_{i}^{2}+\epsilon},$ (10)
where $p_{i}$ represents the predicted probabilities for the actual class labels $g_{i}$, and $\epsilon$ is a small constant introduced to prevent division by zero.

### v2

@@ Abstract
Deep learning has revolutionized medical and biological imaging, particularly in segmentation tasks. However, segmenting biological cells remains challenging due to the high variability and complexity of cell shapes. Addressing this challenge requires high-quality datasets that accurately represent the diverse morphologies found in biological cells. Existing cell segmentation datasets are often limited by their focus on regular and uniform shapes. In this paper, we introduce a novel benchmark dataset of Ntera-2 (NT2) cells, a pluripotent carcinoma cell line, exhibiting diverse morphologies across multiple stages of differentiation, capturing the intricate and heterogeneous cellular structures that complicate segmentation tasks. To address these challenges, we propose an uncertainty-aware deep learning framework for complex cellular morphology segmentation (MorphoSeg) by incorporating sampling of virtual outliers from low-likelihood regions during training. Our comprehensive experimental evaluations against state-of-the-art baselines demonstrate that MorphoSeg significantly enhances segmentation accuracy, achieving up to a 7.74% increase in the Dice Similarity Coefficient (DSC) and a 28.36 reduction in the Hausdorff Distance. These findings highlight the effectiveness of our dataset and methodology in advancing cell segmentation capabilities, especially for complex and variable cell morphologies. The dataset and source code is publicly available at https://github.com/RanchoGoose/MorphoSeg.
@@ 1 Introduction
To address the aforementioned challenges, we propose a novel uncertainty-aware enhancement algorithm based on the SOTA medical segmentation framework TransUNet [1]. Our approach involves sampling virtual outliers from the low-likelihood regions of the class-conditional distributions during the training stage to improve the decision boundaries of the segmented objects. This strategy enhances the robustness of the deep learning models, making them effective solutions in such complex scenarios. By incorporating uncertainties into the training objective, our proposed deep learning model achieves strong segmentation of cells with random shapes. When compared against several benchmarks, we demonstrate significant improvements with our proposed dataset: the mean Dice Similarity Coefficient (DSC) increased from 80.35% to 86.57% and the mean Hausdorff Distance (HD95) decreased from 21.98 to 15.75. These results confirm that our approach not only mitigates data scarcity but also enhances the model’s generalization capabilities across varied cell morphologies. We have provided the full code and pre-trained models with simple inference instructions to facilitate easy implementation and further research by field researchers.
@@
    Development of a Novel Cell Repository: We introduce a pluripotent carcinoma cell repository specifically curated to support the development and validation of advanced biological cell segmentation models. Unlike conventional datasets that predominantly feature cells with regular, round morphologies, our dataset encompasses cells with complex morphologies and transitional stages, presenting a more challenging segmentation task.
@@
    An Innovative Cell Segmentation Algorithm MorphoSeg: We propose MorphoSeg, a novel approach designed to address the challenges posed by our dataset. Our method leverages synthetic outliers from low-likelihood regions in the latent space to significantly enhance the training performance of state-of-the-art cell segmentation models, especially in scenarios with limited data availability. This approach demonstrates superior performance compared to existing methods that often struggle under similar conditions.
@@
    Open-Source Implementation and Comprehensive Evaluation: We release the complete code and detailed data processing procedures as open-source, enabling researchers in the biological field to readily apply our method to raw images and obtain segmentation masks directly. Our extensive quantitative and qualitative evaluations confirm that the proposed model outperforms current techniques.
@@ 2 Related Work
In recent years, deep learning techniques in computer vision have achieved promising performance in medical image segmentation. Encoder-decoder-based CNN structures have been widely applied, with U-Net and its variants significantly advancing the field [2, 12, 18, 19, 20]. Building upon this structure, UNet++ incorporates densely connected links to enhance the fusion of multi-scale features, improving the analysis of medical images across CT, MRI, and EM modalities [21]. Furthermore, nnU-Net, a self-configuring U-Net, has shown remarkable performance across various segmentation tasks with minimal manual parameter tuning [22]. In terms of segmenting cell structures, a framework named Cellpose representing the CNN-based Unet serires performance with the enhancement . The Cellpose model utilizes a two-headed output architecture to predict cell probabilities and spatial organization, effectively handling cells of regular shapes [6].
@@
Semantic segmentation is pivotal for clinical tasks such as disease detection, differential diagnosis, survival prediction, therapy planning, and treatment response assessment. Besides deep learning methods based on CNN structures, attention mechanisms [23] are extensively employed in semantic segmentation to prioritize salient features. The integration of Vision Transformers (ViTs) into medical segmentation networks has enhanced performance by leveraging self-attention mechanisms [24]. Some research focuses on combining the strengths of transformers with traditional CNN architectures, creating hybrid encoder structures. TransUNet combines CNNs for spatial information extraction with transformers for capturing long-range dependencies, featuring a U-Net structure with self-attention mechanisms to ensure comprehensive image analysis [1]. This integration makes TransUNet particularly effective for detailed medical and biological image segmentation, where preserving spatial hierarchies is crucial. Building upon this concept, SwinUNet [25] incorporates Swin Transformer’s shifted-window approach, effectively capturing fine-grained features in small-scale targets. Additionally, ScaleFormer [26] further enhances performance by adopting a more powerful Transformer-based mixer, demonstrating superior capability in handling complex segmentation tasks.
Other research includes employing channel attention to capture boundary-aware features for enhancing polyp segmentation [27]. CellViT replaces the CNN encoder with a transformer block in the U-Net architecture [28]. The recent SAMed model extends the capabilities of the Segment Anything Model (SAM) [29] to medical imaging by employing a Low-rank Adaptation (LoRA) fine-tuning strategy to adapt the large-scale image encoder for medical contexts effectively [3]. This method illustrates the advantage of using extensive pre-trained models for specialized medical segmentation tasks. However, SAM is designed for multi-class segmentation, posing limitations for binary segmentation scenarios.
Regarding on complex cell structure cell segmentation, some recent works have tried to use geometric-aware approaches. Topological consistency is critical in segmentation tasks, especially for reticular images where even slight changes can significantly affect downstream analysis [30]. For curvilinear structures such as vasculature and road networks, a method leveraging discrete Morse theory has been proposed to capture structural uncertainties and identify error-prone connections [31]. A topology-aware loss function that combines skeleton-based weighting with boundary rectification further enhances segmentation accuracy by emphasizing topologically critical regions [32]. In digital pathology, a semi-supervised approach employs a noise-aware topological consistency loss to learn robust representations from unlabeled images, reducing errors in gland and nuclei segmentation [33]. In contrast to these methods, our approach integrates uncertainty estimation directly into the training loss via an outlier synthesis technique, making it compatible with and complementary to existing topological segmentation frameworks.
@@ 2.2 Virtual Outliers Synthesis
Outlier detection plays a crucial role in data mining and statistics, where outliers are defined as abnormalities or deviations from expected patterns [34]. In deep learning, these outliers are often interpreted as anomalies relative to inlier data, and various methods have been developed to identify them [35]. Unlike traditional approaches that focus on outlier detection, recent works in computer vision have explored synthesizing outliers to enhance the accuracy and robustness of deep learning models, particularly in scenarios where datasets are insufficient, such as the medical image classification [36].
Synthetic data has also been employed to quantify model uncertainty, especially for out-of-distribution (OOD) detection. One study leverages synthetic data to better define the boundaries between known and unknown samples [37], which is particularly relevant for detecting unseen cells or cells at transitional stages. OOD detection methods often struggle with the subtle differences among foreground elements in medical images [38].
Synthesis methods have shown promise in various domains, yet their application to medical and biological image segmentation remains relatively underexplored, with few approaches targeting deep learning model training. A recent work combines image synthesis, Gaussian mixture models, and one-class support vector machines to robustly detect hyperintense lesions in cerebral small vessel disease without being limited by pathology or location [39]. Another research introduces a synthetic data generation pipeline, SinGAN-Seg, which leverages a single training image to produce realistic medical images and corresponding masks, thereby enhancing segmentation performance in scenarios with limited data [40]. Additionally, an adversarial confidence learning framework integrates voxel- and region-wise confidence information into supervised segmentation and synthesis models, resulting in improved visual perception and quantitative precision on clinical datasets [41].
In contrast to these methods, our proposed approach dynamically estimates class-conditional Gaussian distributions during training, shaping the uncertainty surface over time with a specially formulated loss. In our cell segmentation task, segmentation targets—specifically NT2 cells at various developmental stages—are treated as OOD data. The demonstrated effectiveness of virtual outlier synthesis in enhancing OOD detection [37] motivates our approach to incorporate synthesized outliers during training to improve segmentation accuracy.
Figure 2: Overview of the MorphoSeg. MorphoSeg employs the TransUNet architecture [1] as an example. The red dotted lines illustrate the virtual outlier sampling process, where feature representations are modeled as class-conditional Gaussians and virtual outliers $v$ are sampled from regions of low likelihood. These outliers contribute to an uncertainty loss for regularization, and the synthetic segmentation map—with an uncertainty estimation branch $\mathcal{L}_{\text{uncertainty}}$—is trained jointly with the mixed segmentation loss $\mathcal{L}_{\text{seg}}$ as defined in Eq. (12).
@@ 3 Methodology
In this section, we detail our proposed approach and outline the overall model architecture. First, Section 3.1 provides background information and an overview of the TransUNet framework for image segmentation. Next, Section 3.2 describes our enhancement of cell segmentation via virtual outlier sampling. Finally, Section 3.3 introduces the uncertainty-aware training objective used in MorphoSeg.
@@ 3.1 TransUNet for Cell Segmentation
Given an image $\bm{\mathrm{x}}\in\mathbb{R}^{H\times W\times C}$ with a spatial resolution of $H\times W$ and $C$ channels, our objective is to predict the corresponding pixel-wise label map of the same dimensions. For a pure Transformer encoder structure, following [24], we first perform tokenization by reshaping the input $\bm{\mathrm{x}}$ into a sequence of flattened 2D patches $\bm{\mathrm{x}}^{i}_{p}\in\mathbb{R}^{P^{2}\cdot C}|i=1,..,N\}$, where each patch is of size $P\times P$ and $N=\frac{HW}{P^{2}}$ is the number of image patches (i.e., the input sequence length). We map the vectorized patches $\bm{\mathrm{x}}_{p}$ into a latent $D$-dimensional embedding space using a trainable linear projection. To encode the patch spatial information, we learn specific position embeddings which are added to the patch embeddings to retain positional information as follows:
@@
The concatenated sequence $\mathbf{T}$ is processed through the Transformer encoder layers to yield the final representations. Then, we employ multi-headed self-attention (MSA) [23] (Eq. 3) and multi-layer perceptron (MLP) blocks (Eq. 4). Layer normalization (LN) is applied before each block (Eq. 5), as described in the following equations:
@@
Due to the limited dimensions of ViT input tokens, input image patch sizes $\frac{H}{P}\times\frac{W}{P}$ are often much smaller than the original image resolution $H\times W$ in high-resolution medical images, leading to a loss of low-level details (e.g., cell shape and boundary). TransUNet addresses this by using a CNN-Transformer hybrid model, where a CNN first extracts a feature map from the input. Patch embedding is then applied to $1\times 1$ patches from the CNN feature map instead of raw images, ensuring better retention of details. We present the hybrid encoder in Figure 2 because it yields better performance, as demonstrated in the comparative analysis of different encoders presented in Table 1 in Section 5.3.
U-Net Structure and Skip Connections. Following TransUNet, we adopt the standard U-Net design with skip connections, replacing naive upsampling with a Cascaded Upsampler (CUP). After reshaping the feature sequence $\bm{\mathrm{z}}_{L}\in\mathbb{R}^{\frac{HW}{P^{2}}\times D}$ into a $\frac{H}{P}\times\frac{W}{P}\times D$ feature map, CUP employs multiple upsampling blocks to progressively restore full resolution. Each block comprises a 2×2$\times$2 × upsampling operator, a $3\times 3$ convolution, and a ReLU activation, enabling effective multi-scale feature aggregation.
Figure 2 illustrates the overall framework of our approach. One thing to notice is that our method is not highly relied on the TransUNet as the outliers sampling processes lies in the latent space and invarient to the model structure, we present the TransUNet here simply as it represents the Transformer based Unet Structure model framework for cell segmentations. The details of how raw high-resolution images are subdivided into small patches to satisfy the input constraints of the vision transformer and to capture fine cellular details presented in Section 4.2.
@@ 3.2 Robust Cell Segmentation via Virtual Outlier Sampling
Our proposed NT2 cell dataset presents two major challenges. First, high-resolution medical images often contain small, densely packed cells that are highly sensitive to pixel-level variations, making accurate human annotation particularly difficult. Second, there is a lack of sufficient training data for the diverse stages of pluripotent carcinoma cells, which limits the model’s ability to generalize across different developmental stages.
To address these issues, we propose a virtual outlier sampling strategy that is applied exclusively during training. Instead of directly sampling outliers from the input image space $\bm{\mathrm{x}}\in\mathbb{R}^{H\times W\times C}$, we synthesize virtual outliers in the latent feature space $\bm{z}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times C^{\prime}}$, where $H^{\prime}$, $W^{\prime}$, and $C^{\prime}$ differ from $H$, $W$, and $C$ due to the spatial and channel transformations introduced by the encoder network.
Operating in the latent space provides several benefits. First, it reduces computational complexity due to the lower dimensionality of latent representations. Second, it enhances training stability and model robustness, as learning to handle synthetic perturbations in the latent space encourages the model to better capture the underlying data distribution. In contrast, perturbing the input space directly may lead to unstable training and degraded performance.
This strategy also enables the incorporation of an uncertainty loss term, $\mathcal{L}_{\text{uncertainty}}$, which guides the model to account for potential outlier regions during training, further improving generalization and segmentation accuracy under limited and imbalanced data conditions.
Specifically, we opt to sample virtual outliers directly from the output logits of the segmentation map, denoted by $s(\mathbf{x}_{i})$, rather than from earlier latent representations. This strategy not only reduces computational load but also minimizes the risk of training instabilities, such as loss explosion, that could arise from discrepancies between artificially added noise and true label distributions.
@@
where $v_{k}\sim\mathcal{N}(\widehat{\mu}_{k},\widehat{\Sigma}_{k})$ represents the sampled virtual outliers for class $k$. These samples are chosen from a sublevel set defined by a small $\epsilon$, ensuring that the outliers are near the decision boundary of the class. With the virtual outliers $v_{k}$ sampled from the predicted segmentation map $s(\mathbf{x}_{i})$, we effectively capture the inherent uncertainty of the segmentation process. In particular, the set of virtual outliers, $\mathcal{V}_{k}$, is integrated into the segmentation pipeline to update $s(\mathbf{x})$, yielding an uncertainty-enhanced output that we denote as $f_{\text{out}}(\mathbf{x};\theta)$. This updated output is then utilized in our uncertainty-aware loss functions in the $\mathcal{L}_{\text{Dice\_out}}$ and $\mathcal{L}_{\text{CE\_out}}$ in Eq. 14 and  13.
We hypothesize that most of the cell pixels are near the boundary of existing cells as shown in Figure 3. Compared to the ground truth masks shown in the second subfigure in Figure 3, the synthesized outliers contain more information about the cells, which are difficult for humans to annotate due to their small size. This strategy aims to enhance the model’s sensitivity to subtle yet critical variations in cell morphology that might indicate different developmental stages or pathological conditions.
@@
3.3 Uncertainty-aware Training Objective
The objective is to develop an uncertainty-aware deep learning model capable of predicting a segmentation map $\hat{S}$ for input cell images. To this end, we introduce a new overall training objective, which utilizes virtual outliers as discussed in Section 3.2. We start by defining the standard Dice loss between the predicted pixel scores $\hat{y}$ and the ground truth targets $y$:
$\mathcal{L}_{\text{Dice}}(f(\mathbf{x};\theta),y)=1-\frac{2\sum_{i}\hat{y}_{i}y_{i}+\epsilon}{\sum_{i}\hat{y}_{i}^{2}+\sum_{i}y_{i}^{2}+\epsilon},$ (10)
where $\hat{y}_{i}$ represents the predicted probabilities for the actual class labels $y_{i}$, and $\epsilon$ is a small constant introduced to prevent division by zero.